# SlowPoke: Understanding and Detecting On-Chip Fail-Slow Failures in Many-Core Systems

### v1

@@
Abstract.
Many-core architectures are essential for high-performance computing, but their performance is undermined by widespread fail-slow failures. Detecting such failures on-chip is challenging, as prior methods from distributed systems are unsuitable due to strict memory limits and their inability to track failures across the hardware topology.
@@ Abstract.
copyright: none
1. Introduction
Many-core architectures, which integrate tens to thousands of processing cores on a single chip or via multi-chiplet integration, are becoming central to modern high-performance computing. Characterized by their massive parallelism and high aggregate throughput, these processors are designed to accelerate today’s most demanding workloads, including deep neural networks (Zheng et al., 2021; Dhingra et al., 2025; Xu et al., 2025a; Dhingra et al., 2024; Yin et al., 2023), large-scale graph analytics (Eyerman et al., 2018; Slota et al., 2015; Cong and Makarychev, 2012), and scientific simulations (Merchant, 2025; Tian et al., 2022; Ahmadzadeh and Sarbazi-Azad, 2023). As a result, many-core architectures are increasingly popular from edge platforms to supercomputers, often delivering significant gains in performance and energy efficiency compared to traditional multi-core CPUs (Ltaief et al., 2023; Kundu et al., 2025; Wang et al., 2023; Wang and Martínez, 2015).
@@ 1. Introduction
However, the scale and complexity of many cores introduce a critical reliability challenge: the fail-slow failure (Gunawi et al., 2018). Unlike failures that result in system crashes (Pillai et al., 2017; Alagappan et al., 2016) or incorrect outputs (Clement et al., 2009), a fail-slow failure allows a hardware component to remain functionally correct but operate at a significantly reduced speed. The causes are both complex and widespread, spanning multiple layers of the system. First, at the device level, physical degradation from effects like aging-induced transistor and interconnect wear can cause timing violations that reduce operational frequency (Agarwal et al., 2007; Bhardwaj et al., 2012; Haghbayan et al., 2020; Wang et al., 2019a). Concurrently, thermal stress introduce dynamic system-wide slowdowns (Zhou et al., 2019). For example, under heavy workloads, extreme temperatures in DRAM can trigger protective states that cut performance by 24% to 52% (Lin and Lin, 2017; Shukla et al., 2019; Pandey et al., 2024). Finally, bottlenecks emerge from shared resource contention and software bugs as cores compete for access to caches, memory, or the on-chip network (Bjerregaard and Mahadevan, 2006; Xiao et al., 2019).
@@
Several approaches exist for detecting fail-slow failures in large-scale parallel and distributed systems. One common strategy is runtime tracing, which collects detailed event logs during execution (Zhai et al., 2010; Khuvis et al., 2018). Another is graph-based analysis, which constructs a dependency graph of the software tasks. By analyzing communication patterns and critical paths within this graph, these methods identify application-level stragglers that are holding up the overall computation (Wang et al., 2022; Gan et al., 2024; Jin et al., 2020). However, these strategies are not a good fit for on-chip detection due to two major challenges. First, runtime tracing is constrained by strict on-chip memory limits. For example, it takes $502\text{\,}\mathrm{KiB}$ and $2\text{\,}\mathrm{MiB}$ of storage to trace the communication and computation of a 16-core processor for just one second (Figure 2), while a single core has typically less than $64\text{\,}\mathrm{KiB}$ of available SRAM. This makes logging impractical and and forces a difficult trade-off between monitoring overhead and accuracy. Second, graph-based analysis struggles with on-chip failure propagation. A slowdown in one hardware component quickly spreads to its neighbors, hiding the true source. Because existing graph-based methods analyze the software tasks and ignore the underlying hardware topology, they often misidentify the source, blaming a component that is merely slowed by a bottleneck elsewhere on the chip.
@@
Figure 2. Challenges for detecting fail-slow failures in many-core processors: (a) On-chip, off-chip memory capacity versus computation and communication tracing log size for tracing DarkNet-19 for 1 second on a $\times4$ core processor. (b) Slowdown propagates across cores along communication dependencies.
@@
Therefore, an effective detection mechanism for many-core systems must be both lightweight to be efficient and hardware-aware to be accurate. A lightweight design must compress data on-the-fly and operate within tight memory limits, while a hardware-aware approach must analyze the chip’s hardware topology to trace a slowdown back to its true origin.
@@
In this paper, we present SlowPoke [^1] The source code of SlowPoke along with failure datasets and traces are available at https://github.com/pku-lemonade/sloth, a fail-slow detection framework with a comprehensive workflow that spans data collection, storage, and analysis to meet these requirements. First, we propose an eBPF-like probe mechanism that enables lightweight and accurate performance data tracing without interfering with program execution. The probes are defined by a configurable tuple, allowing users to precisely specify insertion points. We further design a sketch-based data structure to compress and store the collected traces efficiently. To narrow down the failure candidates, we employ a group-based fail-core detection in combination with underdetermined linear system-based fail-link inference. To capture both software and hardware propagation of fail-slow failures, we reconstruct the communication dependency graph onto the hardware topology, and incorporate temporal information through hierarchical graph. Finally, we introduce FailRank, an iterative ranking algorithm that computes a fail-slow score for each component, thereby identifying the most likely root cause of fail-slow failures. In summary, our contributions are:
  • SL-Compiler: A programmable fail-slow probe mechanism that leverages compiler-based instrumentation. This allows for fine-grained, customizable event monitoring at runtime with minimal execution overhead.
@@
2. Background
2.1. Many-Core Architectures
@@
Many-core architectures integrate cores into a scalable on-chip network to exploit massive parallelism. Figure 3 illustrates the fundamental components of such an architecture. Inter-core communication is enabled by the on-chip network, which can adopt a variety of topologies. Common structures include 2D-mesh, torus, and ring networks (Zheng et al., 2021; Feng et al., 2024), while more complex topologies such as hypercube and dragonfly have also been proposed (Feng et al., 2023). These topologies can support different communication patterns and, in specific scenarios, provide higher bandwidth. In this work, we focus on mesh-based networks, as they are implementation-friendly and are the most widely adopted topology in practice (Taylor et al., 2003; Gratz et al., 2006). Each core is directly connected to a router and consists of three major components: a control unit, a compute unit, and a scratchpad memory. The control unit manages intra-core resource allocation and task scheduling. The compute unit executes application workloads and integrates heterogeneous accelerators such as vector units or tensor units, depending on requirements. The scratchpad memory is a small, high-speed SRAM placed adjacent to the compute unit, typically less than $64\text{\,}\mathrm{KiB}$ in capacity. It is explicitly managed by software or compilers, thereby simplifying hardware coherence logic (Egger et al., 2006; Banakar et al., 2002).
@@
Figure 3. Overview of a many-core architecture with 2D-mesh network and its intra-tile organization.
@@
Based on the hardware organization above, many-core systems often employ a dataflow execution model to maximize concurrency and overlap computation with communication(Chen et al., 2016; Hegde et al., 2021; Cai et al., 2024). In this paradigm, programs are represented as computation graphs, where nodes denote operations and edges represent data dependencies(Dennis and Misunas, 1974; Nikhil et al., 2002). Execution is data-driven—an operation is triggered as soon as its input data becomes available, avoiding the cost of global synchronization. Tasks are mapped onto cores according to data locality and computational workflow, while the scheduler within each core monitors operation readiness, fetches operands from local or neighboring cores, and dispatches them to compute units. Intermediate results are routed through the NoC according to the communication pattern derived from the computation graph. This fine-grained pipelining of data movement and processing improves utilization, but also creates tight coupling between cores: delays in one core’s execution can propagate through dependent operations across the chip, amplifying the performance impact of fail-slow failures.
2.2. Hardware Failures
@@
In many-core systems, hardware components are subject to a wide range of failures, which may manifest in different ways but invariably influence normal program execution. Based on this behavior, common hardware failures can be broadly classified into two categories: fail-slow and fail-stop(Gunawi et al., 2018; Panda et al., 2019).
@@
For cores, bit flips can alter register values, leading to incorrect computation results or even program crashes (Sastry Hari et al., 2009; He et al., 2023), which are characteristic of fail-stop failures. In contrast, performance variations caused by internal resource contention, workload imbalance, or hardware aging typically degrade overall performance (Tang et al., 2018; Zheng et al., 2022b), and are categorized as fail-slow failures. For memory, permanent physical defects as well as software bugs or mismanagement can trigger system crashes, thus are fail-stop failures. Meanwhile, memory leaks or buffer overflows could consume significant system resources, causing response delays and performance degradation, which belong to fail-slow failures(Lu et al., 2022; Xu et al., 2025b). For NoC, transient bandwidth reductions on links or low router port utilization can decrease overall throughput and prolong program execution time(Chang et al., 2011; Tsoutsouras et al., 2017), both of which are fail-slow failures. In contrast, routing-induced deadlocks can completely stall communication, making them fail-stop failures(Taheri et al., 2022).
@@
2.3. Performance Variance Modeling
Beside fail-slow failures, hardware performance has inherent variability(Marathe et al., 2017). To accurately capture the communication costs in many-core systems, we model those key performance parameters as random variables. Specifically, we model two main contributors to operation latency: computing capacity for cores and bandwidth for NoC.
Computing Capacity of each core, which affected by thermal pressure, throttling mechanism and so on, is assumed to follow a normal distribution (Inadomi et al., 2015; Romanescu et al., 2006). This distribution reflects the aggregated effect of microarchitectural factors such as clock frequency variation, temporary resource contention, and thermal-induced fluctuations, all of which tend to produce near-symmetric deviation around a nominal value. We denote the computing capacity $P_{core}\sim N(\mu_{c},\sigma^{2}_{c})$ where $\mu_{c}$ represents the expected processing capacity and $\sigma^{2}_{c}$ quantifies the runtime variability.
Bandwidth Fluctuation. The latency associated with on-chip interconnect links is subject to more asymmetric and bursty behavior, particularly under dynamic network load or in the presence of degraded hardware(Hernández et al., 2010). To model this, we adopt a Gamma distribution (Marathe et al., 2017), $T_{link}\sim Gamma(\alpha,\beta)$, which provides flexibility in capturing latency variation caused by bandwidth. The shape parameter controls the burstiness of communication delays, while the rate parameter reflects average link throughput under contention. This distribution is particularly effective in approximating observed behaviors such as occasional long-tailed delays due to arbitration, flit contention, or rerouting in degraded paths.
Figure 4. Overview of SlowPoke.
3. The SlowPoke Framework
@@
SlowPoke is an automated framework for detecting and locating fail-slow failures in many-core systems. As shown in Figure 4, the framework is composed of three main components. SL-Compiler instruments the workload with lightweight probes, collecting runtime execution trace with minimal overhead. SL-Recorder then filters unrelevant trace and compresses redundant ones, producing a compact yet representative communication and execution patterns list. Finally, SL-Tracer leverages the compressed data to construct multi-level communication graph and perform FailRank algorithm on it, pinpointing fail-slow cores or links responsible for performance degradation. Together, these components form a systematic pipeline that bridges low-overhead monitoring with high-accuracy failures diagnosis.
3.1. Input and Output
@@
As shown in the left of Figure 4, the inputs to SlowPoke consist of: (1) Many-Core Workload: workloads are composed of basic I/O, communication, and compute instructions such as read, recv, and conv. The framework supports arbitrary task-to-core mappings (Zheng et al., 2022a; Kwon et al., 2020; Hanson et al., 2023). In our implementation, we use a Gemini-like (Cai et al., 2024) mapping for evaluation; (2) Arch Config: the specification of target hardware architecture, which enables analysis across different many-core systems; (3) Probe Config: parameters that determine what data to collect, where to place probes, and how to store the collected traces data; (4) Failure Model: definitions of fail-slow behavior, including duration, degradation rate, and other settings.
@@
Given these inputs, SlowPoke automatically generates probed instructions, collects runtime traces, and performs fail-slow analysis. The outputs fall into two categories: fail-slow detection results, including fail-slow root cause, SL-Recorder’s optimal parameters, and the overhead of data collection and storage; and system-level information, such as total simulation time and execution latency.
3.2. SL-Compiler
This SL-Compiler takes workload mapping as input and is responsible for instrumenting program code with lightweight performance probes before execution. Inspired by the modular and event-driven design of eBPF (extended Berkeley Packet Filter), we decompose the monitoring process into three key abstractions. Probe fragments define what data to collect, insertion tracepoints specify where to place the probes, and performance data maps determine how to store the collected data. Figure 5 shows an example of SL-Compiler probing. The probed compute graph has detailed monitoring fragments without changing the program’s logic or significantly affecting its runtime performance.
Figure 5. Examples of SL-Compiler probing.
The probe strategy is defined by a five-tuple. As summarized in Table 1, each abstraction is parameterized by several fine-grained components. Fragment controls the probe fragment; Type and Location jointly specify the insertion tracepoint; Level and Structure define the data map for managing collected traces. By adapting the core ideas of eBPF, we create a user-friendly and efficient monitoring mechanism tailored for many-core architectures.
Probe fragment. To collect performance data without disrupting program execution, we introduce the probe fragment, a lightweight pseudo-instruction. As shown in Table 1, fragments can capture computation-related events (Exec), communication activities (Route), or memory accesses (Mem). Figure 5b illustrates this case: by configuring Fragment=Exec, probes are injected on compute instructions to record execution time. Similarly, Figure 5c shows the effect of setting Fragment=Route, which inserts probes on communication instructions to measure transmission latency. Crucially, these fragments do not interfere with the main compute pipeline, making them a portable tool for performance monitoring across different many-core systems.
Insertion tracepoint. To precisely control where fragments are placed, SL-Compiler introduces configurable insertion tracepoints, defined by Type (instruction filter) and Location (relative position), allowing flexible combination. As shown in Table 1, Type can be Comp, Comm, or IO to filter the target instruction type, while Location specifies whether the probe is anchored before, after, or surrounding the instruction. For example, by combining Type=Comp and Location=Post, Figure 5b insert probes after every compute instruction. These tracepoints are automatically inserted by SL-Compiler based on the static workload mapping, enabling complete monitoring coverage without manual work. Therefore, users can easily switch to different monitoring strategies to meet their needs.
Table 1. Components of probe’s five-tuple definition.

### v2

# Towards Efficient and Accurate Detection of On-Chip Fail-Slow Failures for Many-Core Accelerators

@@
Abstract
Many-core accelerators are essential for high-performance deep learning, but their performance is undermined by widespread fail-slow failures. Detecting such failures on-chip is challenging, as prior methods from distributed systems are unsuitable due to strict memory limits and their inability to track failures across the hardware topology.
@@ Abstract
I Introduction
Many-core accelerators, integrating tens to thousands of processing cores on a single chip or via multi-chiplet packaging, have emerged as the backbone of modern applications. To support massive parallelism, these architectures rely on a tightly coupled design, where cores are interconnected via a high-bandwidth Network-on-Chip (NoC) to enable frequent, low-latency communication. This fabric allows them to efficiently handle throughput-demanding workloads [83, 14, 78, 13, 80], ranging from large-scale model training and inference to graph analytics [17, 64, 11] and scientific simulations [50, 69, 2]. As a result, many-core accelerators are increasingly deployed across diverse platforms from edge devices to data centers, delivering significant gains in performance and energy efficiency compared to traditional multi-core CPUs and GPGPUs [44, 35, 73, 74].
@@ I Introduction
The characteristics of fail-slow failures, especially the transiency, mandate an on-chip detection paradigm, as traditional off-chip mechanisms are fundamentally ill-suited to capture such phenomena. On the one hand, the transient nature of these failures creates a critical detection gap for external observers. Off-chip monitoring typically relies on periodic polling or remote telemetry, operating at a coarse granularity and with significant transmission latency. By the time the status data reaches an external host, the micro-architectural stall effectively vanishes, rendering the external detector blind to the fleeting root cause. On the other hand, precise diagnosis requires high-resolution traffic data, which inevitably collides with the I/O bandwidth wall. Streaming continuous, fine-grained execution traces off-chip imposes a prohibitive bandwidth overhead. This heavy data movement not only effectively saturates limited I/O resources but also induces a parasitic effect: the diagnosis traffic competes with the actual workload for memory access, distorting the very system behavior it attempts to measure. Therefore, effective identification necessitates a shift to on-chip detection. Crucially, to accommodate hardware constraints, this approach must be lightweight and capable of performing in-situ statistical aggregation, capturing the failure root cause locally without the latency or bandwidth penalties of external analysis.
TABLE I: Comparison of hardware failure detection approaches in different scenes.
Target Scenario Detection Approach Failure Pattern Failure Model Validation Method Memory Overhead
Cloud System Sieve[15] Fail-Slow Latency Simulation 8 GiB
Micro Service Groot[71] Fail-Stop Fail-Slow Error Degradation Real Plat. 20 GiB
Sleuth[21] Fail-Slow Contention Simulation Real Plat. 84 MiB
Sage[20] Fail-Slow Latency Contention Real Plat. Not Mentioned
Distributed System IASO[55] Fail-Slow Timeout Real Plat. 1.6 MiB Per Hour
Perseus[46] Fail-Slow Latency Real Plat. Not Mentioned
ADR[45] Fail-Slow Packet Loss Latency Real Plat. Not Mentioned
Many-Core Accelerator SLOTH (Ours) Fail-Slow Degradation Simulation 150 KiB
While fail-slow detection has been explored in microservices and distributed systems (Table I), adapting these paradigms to many-core accelerators is impeded by two challenges. First, many-core systems have strict on-chip memory limits. Standard runtime tracing strategies [55, 46] and graph-based analyses [71, 21, 33] rely on capturing extensive execution logs, incurring massive storage overheads (e.g., Gigabytes for Sieve and Groot) that contradict the hardware reality shown in Figure 2a. Even a minimal 1-second trace of DarkNet-19 necessitates $2\text{\,}\mathrm{MiB}$ just for communication logs; this creates an unbridgeable gap with on-chip SRAM capacity (typically $64\text{\,}\mathrm{KiB}$ per core), rendering continuous in-situ tracing spatially infeasible. Second, existing methods are lack of hardware topological awareness. while graph-based methods analyze software dependencies, they struggle to distinguish root causes from propagated symptoms due to the complex on-chip interconnect topology. As illustrated in Figure 2b, a fail-slow failure in Core 1 triggers a cascading slowdown: the delayed data transmission propagates along the communication links, forcing downstream cores (Core 2 and Core 3) into involuntary wait states. Because existing methods focus on application-level stragglers without accounting for the underlying hardware backpressure dynamics, they inevitably conflate the symptoms with the cause, misidentifying the backpressured victims as the culprits.
@@
Figure 2: Challenges for detecting fail-slow failures: (a) On-chip, off-chip memory capacity versus computation and communication tracing log size for tracing DarkNet-19 for 1 second on a $\times4$ core processor. (b) Slowdown propagates across cores along communication dependencies.
@@
Therefore, an effective detection mechanism for many-core accelerator must be both lightweight to be efficient and hardware-aware to be accurate. A lightweight design must operate within tight memory limits, while a hardware-aware approach must analyze the accelerator’s hardware topology to localize the root cause.
In this paper, we present SLOTH [^1] The source code of SLOTH along with failure datasets and traces are available at https://anonymous.4open.science/r/sloth-91D7(Slowdown Localization via On-chip Tracing and Hierarchical graph analysis), a fail-slow detection framework with a comprehensive workflow that spans data collection, compression, storage, and analysis to meet these requirements.
@@
First, we propose a workload-aware probe mechanism that enables lightweight performance data tracing without interfering with program execution. It automatically analyzes the computation graph to identify insertion points, collecting operator-level performance metrics. We further design a sketch-based data structure to compress and store the collected traces efficiently. To narrow down the failure candidates, we employ a group-based fail-core detection in combination with underdetermined linear system-based fail-link inference. To capture both software and hardware propagation of fail-slow failures, we reconstruct the communication dependency graph onto the hardware topology, and incorporate temporal information through hierarchical extension. Finally, we introduce FailRank, an iterative ranking algorithm that computes a fail-slow score for each component, thereby identifying the most likely root cause of fail-slow failures. In summary, our contributions are:
  • SL-Compiler: A workload-aware probe mechanism that leverages compiler-based instrumentation to automatically insert probe at operator boundaries. This enables operator-level tracing with minimal overhead while capturing fine-grained metrics.
@@
II Background
II-A Many-Core DNN Accelerators
@@
Many-core DNN accelerators integrate numerous cores into a scalable on-chip network to exploit massive parallelism for deep learning workloads. Figure 3 illustrates the fundamental components of such an architecture. Inter-core communication is enabled by the Network-on-Chip (NoC), which can adopt various topologies. Common structures include 2D-mesh, torus, and ring networks [83, 18], while more complex topologies such as hypercube and dragonfly have also been proposed [19]. These topologies can support different communication patterns and, in specific scenarios, provide higher bandwidth. In this work, we focus on mesh-based networks, as they are implementation-friendly and are the most widely adopted topology in practice [68, 22]. Each core is directly connected to a router and consists of three major components: a control unit, a compute unit, and a scratchpad memory. The control unit manages intra-core resource allocation and task scheduling. The compute unit executes DNN operators and typically integrates specialized accelerators such as vector units or tensor units optimized for matrix operations. The scratchpad memory is a small, high-speed SRAM placed adjacent to the compute unit, typically less than $64\text{\,}\mathrm{KiB}$ in capacity, storing activations, weights, and intermediate results. It is explicitly managed by software or compilers, thereby simplifying hardware coherence logic [16, 4].
@@
Figure 3: Overview of a many-core DNN accelerator with 2D-mesh network and its intra-tile organization.
@@
Based on the hardware organization above, many-core DNN accelerators employ a dataflow execution model to maximize concurrency and overlap computation with communication[9, 28, 7]. In this paradigm, DNN models are represented as computation graphs, where nodes denote operators (e.g., convolution, pooling) and edges represent data dependencies between layers[12, 54]. Execution is data-driven—an operator is triggered as soon as its input features and weights become available, avoiding the cost of global synchronization. DNN operators are mapped onto cores according to data locality and the layer-wise computational workflow, while the scheduler within each core monitors operator readiness, fetches operands from local or neighboring cores, and dispatches them to compute units. Intermediate features are routed through the NoC according to the communication pattern derived from the DNN computation graph. This fine-grained pipelining of data movement and operator processing improves hardware utilization, but also creates tight coupling between cores: delays in one core’s execution can propagate through dependent operators across the chip, amplifying the performance impact of fail-slow failures in DNN inference.
II-B Hardware Failures
@@
In many-core DNN accelerators, hardware components are subject to a wide range of failures that invariably degrade inference performance. Based on their behavioral characteristics, common hardware failures are classified into two categories: fail-slow and fail-stop[23, 55].
@@
For cores, bit flips can alter register values, leading to incorrect computation results or even program crashes [60, 27], which are characteristic of fail-stop failures. In contrast, performance variations caused by internal resource contention, workload imbalance, or hardware aging typically degrade overall performance [67, 85], and are categorized as fail-slow failures. For memory, permanent physical defects as well as software bugs or mismanagement can trigger system crashes, thus are fail-stop failures. Meanwhile, memory leaks or buffer overflows could consume significant system resources, causing response delays and performance degradation, which belong to fail-slow failures[47, 79]. For NoC, transient bandwidth reductions on links or low router port utilization can decrease overall throughput and prolong program execution time[8, 70], both of which are fail-slow failures. In contrast, routing-induced deadlocks can completely stall communication, making them fail-stop failures[66].
@@
Figure 4: Overview of SLOTH.
III The SLOTH Framework
@@
SLOTH is an automated framework for detecting and locating fail-slow failures in many-core DNN accelerators. As shown in Figure 4, the framework is composed of three main components. SL-Compiler instruments DNN operators with lightweight probes at operator boundaries, collecting fine-grained runtime execution traces with minimal overhead. SL-Recorder then filters irrelevant traces and compresses redundant ones, producing a compact yet representative list of communication and execution patterns. Finally, SL-Tracer leverages the compressed data to construct multi-level communication graphs and performs the FailRank algorithm on them, pinpointing fail-slow cores or links responsible for performance degradation. Together, these components form a systematic pipeline that bridges low-overhead monitoring with high-accuracy failure diagnosis.
III-A Input and Output
@@
As shown in the left of Figure 4, the inputs to SLOTH consist of: (1) DNN Workload: DNN models are composed of operators such as convolution, pooling, and data movement instructions including read, send, and recv. The framework supports arbitrary operator-to-core mappings [84, 36, 25]. In our implementation, we use a Gemini-like [7] mapping strategy for evaluation; (2) Arch Config: the specification of the target many-core DNN accelerator architecture, which enables analysis across different hardware platforms; (3) Probe Config: parameters that determine what execution information to collect, where to place probes at operator boundaries, and how to store the collected trace data; (4) Failure Model: definitions of fail-slow behavior, including failure duration, performance degradation rate, and affected components.
@@
The outputs fall into two categories: fail-slow detection results, including the root causes, SL-Recorder’s optimal compression parameters, and the overhead of data collection and storage; and system-level metrics, such as total inference time and per-layer execution latency.
III-B SL-Compiler
Traditional tracing approaches face a trade-off between coverage and overhead: exhaustive collection is prohibitively expensive, while reactive collection risks missing critical failure signatures. To address this challenge, we propose SL-Compiler, a DNN-aware probe instrumentation tool that analyzes the hierarchical structure and explicit data dependencies of DNN workloads to insert eBPF (extended Berkeley Packet Filter) inspired probes, reducing monitoring overhead while preserving the ability to capture fail-slow behaviors.
TABLE II: Components of probe’s five-tuple definition.